OpenStudy (anonymous):

The square root of x-10 = 4

OpenStudy (anonymous):

\[\sqrt{x-10}=4\]

OpenStudy (anonymous):

is that what it is supposed to be?

OpenStudy (abb0t):

No. Why did you only square one side?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

im trying to figure out the original equation @abb0t

OpenStudy (anonymous):

It's supposed to be the first one

OpenStudy (abb0t):

What are you being asked to do here @Mady451

OpenStudy (anonymous):

ok, so square both sides and you get x-10=16

OpenStudy (anonymous):

x=26

OpenStudy (jhannybean):

^ don't give answers.

OpenStudy (anonymous):

x-10=16 and then after that you add ten to each side

OpenStudy (anonymous):

which would cancel out the 10 on the left and make the 16, 26

OpenStudy (anonymous):

Thanks